# LLMs for Generalizable Language-Conditioned Policy Learning under Minimal Data Requirements

### v1

@@ A.4 Testing Environments
This paper uses the Minigrid-BabyAI environment to benchmark its method. This choice was motivated by several factors. Most importantly, we require a sandbox environment in which a wide range of goal reaching tasks can be expressed in natural language. Robotic environments (Todorov et al. (2012); James et al. (2019)) were excluded due to precise control of robotic components being beyond LLM’s prior knowledge and the need to discretize continuous actions to match LLM’s tokenized output. Additionally, 3D environments (Fan et al. (2022); Puig et al. (2018)) were not considered due to computational constraints. Text-based games (Côté et al. (2018); Shridhar et al. (2021b)) were also excluded as they involve high-level text interactions, contrary to this paper’s focus on low-level control task for language models.
@@
Given the significant computational resources and time required to perform all three steps of our pipeline, in particular fine-tuning an LLM agent, our current scope is necessarily limited to BabyAI. Nonetheless, the insights derived from this controlled setting are broadly applicable and provide a foundation for future work in environments with similar tabular structures, such as NetHack (Küttler et al. (2020)) and Overcooked (Carroll et al. (2020)), which differ mainly in thematic focus (video game dungeon crawling and collaborative cooking, respectively).
Appendix B Experimental details
B.1 Data Collection
@@
To collect the data ${\mathcal{D}}$ used throughout the experiments, we rely on the default goal-oriented policies from the BabyAI environment. We denote these policies by $\pi^{\beta}(\cdot;g)$. Our data collection policy that is random mixture of the policies $\pi^{\beta}(\cdot;g)$. Given a randomly sampled initial state $x_{0}\in{\mathcal{X}}$ and an unknown goal $g$ randomly sampled from the set of original BabyAI language commands, we let the agent interact with the environment according to $\pi^{\beta}(\cdot;g)$ until either $g$ is reached or the limit of 500 steps is reached.This policy simulates agents attempting to accomplish multiple task within the environment. Examples of real-world unlabeled data that could be generated from such policy include CCTV footage of employees at work, logs of medical procedures performed on a patient, or YouTube videos.
@@
Refer to Appendix C.1 for an analysis of how the data collection policy affects our pipeline, comparing goal-oriented data collection with fully random data collection.
B.2 Abstraction Function
@@
The abstraction function utilizes contextual understanding of LLMs to identify goal-relevant features. We prompt an LLM with the given goal, a randomly sampled state representation as an exmple, and two in-context examples of the expected output. Figure B.1 shows the prompt template used. The LLM returns the goal relevant features which are then passed to a python function that processes states according to the following rules:
@@
For this step we use the Llama-3-70B-Instruct language model with the following parameters: {temperature: 0, top k: 1, maximum number of tokens: 8000}.
@@
Figure B.1: Prompt template for selecting the relevant features to achieve the goal.
@@ 27. The agent is facing up
Figure B.2: An example of BabyAI textualized state before state abstraction.
@@
Figure B.3: Textualized state from B.2 after applying state abstraction for the goal “pick up a ball”.
B.3 Reward Shaping
@@
In the first stage, a large language model LLM, Llama-3-70B-Instruct, is utilized to generate a supervised dataset of labeled goals. The LLM is configured with parameters {temperature: 0, top-k: 1, max tokens: 8000}, using the prompt template shown in Figure B.4. For each goal $g$, up to 5000 states are randomly sampled from $\tilde{\mathcal{S}}^{g}$ and labeled.
In the second stage, a collection of neural networks is trained on this dataset. The state representations are transformed from text to a grid format. The network architecture consists of a small convolutional neural network with one convolutional layer (output dimension: 32, kernel size: (2,2)), followed by two linear layers (hidden dimension: 32, output dimension: 1). A Sigmoid activation function is applied after the final linear layer, and ReLU is used after all other layers. Dropout layers are added before each linear layer. The network is trained with the following hyperparameters: learning rate of 1e-5, maximum of 3000 epochs, and dropout rate of 0.1. The dataset is split into training and validation sets (90%/10%), and the model weights with the lowest validation loss are retained.
@@
Figure B.4: Prompt template for labeling states as goal states or not.
B.4 Tabular Q-learning
@@
$Q_{new}^{g}[s_{t},a_{t}]\leftarrow(1-\alpha)Q^{g}[s_{t},a_{t}]+\alpha\left(R[s_{t},a_{t}]+\gamma\max_{a}Q^{g}[s_{t+1},a]\right).$
The learning rate $\alpha$ is set to 0.1, and the discount factor $\gamma$ is set to 0.7. Subsequent states $s_{t+1}$ are restricted to transitions observed in $\mathcal{D}$. Iterations stop when $||Q_{new}^{g}-Q^{g}||_{\infty}<\epsilon$, where $\epsilon=1\times 10^{-6}$.
B.5 LLM Fine-tuning
Table B.1: Fine-tuning hyperparameters
{tblr}
width = colspec = Q[713]Q[169], row1 = c, cell22 = c, cell32 = c, cell42 = c, cell52 = c, cell62 = c, cell72 = c, cell82 = c, cell92 = c, cell102 = c, cell112 = c, hline1,12 = -0.08em, hline2 = -, Hyperparameter & Value
Batch size (per device) 10
Learning rate 2e-5
Maximum Gradient norm 0.3
Warmup ratio 0.01
Maximum number of epochs 3
LORA rank 512
LORA alpha 512
LORA dropout 0.1
Split train/val ratio 0.1
Tensor type bf16
TEDUO’s step 3 involves fine-tuning a large language model using the generated supervised dataset ${\mathcal{D}}^{SFT}$. In this paper, the fine-tuned model is Llama-3-8B-Instruct. We use Low-Rank Adaptation (Hu et al. (2021)) to reduce the compute cost. The hyperparameters used for the fine-tuning step are detailed in Table B.1. The model weights with the lowest validation loss are retained. The fine-tunings have been realised on a cluster of 4 A100 (80GB VRAM). The computing power provided in figure 4 is determined by multiplying the number of GPU hours by the peak Tflops (312 for A100 in bf16) and the estimated utilisation rate (90%).
@@ B.5 LLM Fine-tuning
Figure B.5: This prompt template is employed to generate a sequence of optimal actions to achieve the given goal while being in the given state.
B.6 Baselines
@@
LLMs (vanilla). The vanilla Large Language Model baseline utilizes Llama-3-8B-Instruct or Llama-3-70B-Instruct prompted with the template shown in Figure B.5. This prompt provides basic information about the environment, current goal, and a textual (non abstracted) representation of the state.
LLMs (in-context + CoT). This baseline extends the vanilla LLM approach by using the prompt in Figure B.6, which includes detailed environment information, similar to a game manual, as described in (Wu et al., 2023). It also integrates expert demonstrations using textual grid examples and goals with their optimal action sequences. The “Chain-of-Thought” (CoT) prompting technique is employed to guide the LLM through multi-step reasoning and self-reflection.
Figure B.6: This prompt template is employed to generate a sequence of optimal actions to achieve the given goal while being in the given state. It uses in-context learning and Chain-of-Thought prompting.
B.7 Evaluation Setup
The evaluation setup outlined in Table 1 includes training and testing configurations for both environments and goals.
Goals. Training goals are defined as the goal contained in $\mathcal{G}^{tr}$, a subset of natural language instructions provided by BabyAI without any modifications. Testing goals differ both grammatically and semantically from training goals. They are derived from BabyAI’s original instructions, distinct from $\mathcal{G}^{tr}$, and reformulated using alternative phrasings and synonyms. Tables B.2, B.3, and B.4 provide the alternative formulations and synonyms for objects and colors used in these reformulations.
Table B.2: Alternative formulations for the natural language commands.
@@ B.7 Evaluation Setup
Table B.3: Synonyms used for the objects.
@@
Table B.4: Synonyms used for the colors.
@@
Appendix C Additional Results
C.1 Data collection policy
To empirically analyze this, we consider two randomized policies:
  • B) Random policy: Actions are drawn uniformly at random from the action space. This policy represents an agent that explores the environment without a specific goal. Although this scenario is less common in real-world settings—where agents typically pursue objectives—it remains applicable forSo batch RL, particularly when learning from untrained agents with no prior knowledge.
Figure C.7: Impact of data collection policy. The x-axis shows the proportion of data ${\mathcal{D}}$ collected with policy A vs. policy B for a fixed size of ${\mathcal{D}}$. a) The y-axis shows $|\mathcal{S}_{0}^{g}|:=\{s_{0}^{g}:(g,s_{0}^{g},[a_{0}^{g,*},a_{1}^{g,*},...])\in{\mathcal{D}}^{SFT}\}$, i.e. the number of unique initial abstract states $s_{0}^{g}$ for which $g$ is reachable with the learned policy $\pi^{g}$. Values are min-max normalized across all 5 mixture policies. b) The y-axis shows the same values as plot a), averaged across 14 goals, bars represent the standard error.
Figure C.7 illustrates that the optimal data collection policy varies by goal. For some goals policy A works better, while for others it is the fully random policy. Importantly, a comparable amount of synthetic action sequences for fine-tunning the LLM can be extracted using either policy A or B. Averaging across all goals, we find that policy A tends to perform better for goals in ${\mathcal{G}}_{\pi}$ than those not in ${\mathcal{G}}_{\pi}$. Future work could explore optimizing the set of training goals $\mathcal{G}^{tr}$ to maximize the alignment of $\mathcal{G}^{tr}$ with a given dataset $\mathcal{D}$. Yet, the necessity of aligning ${\mathcal{D}}$ and $\mathcal{G}^{tr}$ is moderated by two factors. First, as shown in subsection 5.4, the abstraction function reduces the complexity of the abstract MDPs, requiring fewer data samples. Second, since extending the list of goals in ${\mathcal{G}}^{tr}$ is computationally inexpensive, we can continually seek better alignment.
C.2 Abstraction function
Figure C.8: Reduction in count of unique states due to applying the $\mathrm{LLM}(g)$ abstraction functions and the relative size of the reduced abstract feature space ${\mathcal{S}}^{g}_{\phi}$, containing only features necessary to identify the completion of a goal $g$.
Appendix D Reward shaping evaluation
Table D.5: Reward Shaping Benchmark. The accuracy, precision, and recall metrics are computed with a classification threshold ensuring at least 95% precision.
{tblr}
Go to a box
0.90 89 96 38
Pick up a ball 0.75 98 95 83
Open a door 0.92 85 95 85
Go to red door 0.98 94 100 0.2
Go to the tile (5,6) 1.0 100 100 100
Put a box next to a blue ball 0.64 100 100 25
Table D.5 shows the performance of $R_{\theta}(\cdot;g)$ for various types of goals compared to ground truth rewards. The benchmark setup is consistent with the main experiments; details are provided in the Appendix B.3. All goals achieve 95% precision, a crucial metric since false positives lead to generating incorrect data points for ${\mathcal{D}}^{SFT}$ in TEDUO’s step 2. Conversely, false negatives only reduce data points in ${\mathcal{D}}^{SFT}$, which is less critical given our synthetic data abundance (see Section 5.4). Performance varies across goals; for instance, ”go to the red door” has low recall (0.2%), likely due to limited positive examples in the dataset. Expanding the dataset could improve such outcomes.
D.1 Benchmark results per goal category
Table D.6: Online evaluation of generalization performance split per goal category. This is the success rate [%] presented in Table 1 with the 400 $(g,s_{0}^{g})$ grouped by goal category. Standard error in brackets.
{tblr}
TEDUO: steps 1 2 + BabyAI-IL-bot train train 46 (± 2.0) 92 (± 1.2) 100 (± 0.0) 7 (± 2.9)
test train 30 (± 1.6) 58 (± 1.7) 100 (± 0.0) 6 (± 2.1)
train test 5 (± 1.2) 44 (± 2.5) 4 (± 0.9) 0 (± 0.0)
test test 7 (± 1.2) 40 (± 2.2) 5 (± 0.9) 0 (± 0.0)
TEDUO (Llama-3-8B) train train 46 (± 2.3) 85 (± 1.3) 100 (± 0.0) 0 (± 0.0)
test train 39 (± 2.4) 65 (± 2.0) 100 (± 0.0) 0 (± 0.0)
train test 20 (± 3.8) 87 (± 3.2) 83 (± 1.8) 0 (± 0.0)
test test 26 (± 3.0) 70 (± 2.8) 61 (± 2.6) 0 (± 0.0)

### v2

@@ C.1 Choice of Environments
BabyAI. This paper primarily uses the Minigrid-BabyAI environment to benchmark its method. This choice was motivated by several factors. Most importantly, we require a sandbox environment in which a wide range of goal reaching tasks can be expressed in natural language. Robotic environments ((Todorov et al., 2012; James et al., 2019)) were excluded due to precise control of robotic components being beyond LLM’s prior knowledge and the need to discretize continuous actions to match LLM’s tokenized output. Additionally, 3D environments ((Fan et al., 2022; Puig et al., 2018)) were not considered due to computational constraints. Text-based games ((Côté et al., 2018; Shridhar et al., 2021b)) were also excluded as they involve high-level text interactions, contrary to this paper’s focus on low-level control task for language models.
@@
Given the significant computational resources and time required to perform all three steps of our pipeline, in particular fine-tuning an LLM agent, our current scope is necessarily limited to BabyAI. Nonetheless, the insights derived from this controlled setting are broadly applicable and provide a foundation for future work in environments with similar tabular structures, such as NetHack ((Küttler et al., 2020)) and Overcooked ((Carroll et al., 2020)), which differ mainly in thematic focus (video game dungeon crawling and collaborative cooking, respectively).
Webshop. To demonstrate the ability of TEDUO to scale beyond grid-world environments, this section proposes an additional benchmark of TEDUO in the Webshop environment (Yao et al., 2023a). The development of digital agents which are can excel at tasks performed on computers, such as web navigation, coding, and operating office software is one of the most promising applications emerging from the LLM-Reinforcement Learning synergy. Webshop is a simulated e-commerce environment where an agent given product requirements must locate corresponding items by navigating a website. This environment is particularly relevant due to its dynamically evolving action space, presenting two key challenges:
Dynamic UI Actions: The available actions vary by state due to changes in the website’s UI elements. We address this challenge by concatenating the available actions with the state description. Linguistic Action Requirements: The agent must generate linguistic inputs (e.g., search queries) to use the search bar. This highlights the necessity for the fine-tuned LLM to avoid catastrophic forgetting (Luo et al., 2025) and produce coherent keywords for narrowing searches.
C.2 Guide for Practitioners
This section provides a high-level summary of the adaptations required to deploy TEDUO in a new environment. The key modifications involve adjusting the stages where domain knowledge is incorporated. Specifically, the following prompts must be adapted:
  1. 1.
    The prompt for generating the abstraction function (Appendix C.4).
  2. 2.
    The prompt for reward labeling (Appendix C.5).
  3. 3.
    The prompt used at inference time (Appendix C.7).
At these stages, prior knowledge about the environment’s dynamics, the semantics of instructions, or any contextual information that can facilitate goal-conditioned policy learning in Step 3 can be integrated via natural language. The provided prompt examples illustrate how such adaptations can be implemented for two different environments.
Additionally, the offline Reinforcement Learning method used to solve the abstract Markov Decision Processes in Step 2 should be tailored to the characteristics of the environment. This includes considering factors such as modality, the scale of the state and action spaces, and the number of initial states.
C.3 Data Collection
As mentioned in the main body of this paper, our pipeline makes no assumption on the data collection policy used.
C.3.1 BabyAI adaptation
@@
In the experiments with the BabyAI environment, we rely on the default goal-oriented policies from the BabyAI environment. We denote these policies by $\pi^{\beta}(\cdot;g)$. Our data collection policy is then a random mixture of the policies $\pi^{\beta}(\cdot;g)$. Given a randomly sampled initial state $x_{0}\in{\mathcal{X}}$ and an unknown goal $g$ randomly sampled from the set of original BabyAI language commands, we let the agent interact with the environment according to $\pi^{\beta}(\cdot;g)$ until either $g$ is reached or the limit of 500 steps is reached. This policy simulates agents attempting to accomplish multiple task within the environment. Examples of real-world unlabeled data that could be generated from such policy include CCTV footage of employees at work, logs of medical procedures performed on a patient, or YouTube videos.
@@
Refer to Appendix B.2 for an analysis of how the data collection policy affects our pipeline, comparing goal-oriented data collection with fully random data collection.
C.3.2 Webshop adaptation
In the experiment with the Webshop environment, to demonstrate that our pipeline can learn from non-expert demonstrations, we design a goal-conditioned data collection policy, $\pi^{\beta}(\cdot;g)$, which relies on a pre-trained LLM and random actions. When encountering a ”search” webpage that allows keyword input for locating the target object, $\pi^{\beta}(\cdot;g)$ prompts a pre-trained LLM with instruction $g$ to generate relevant keywords. In all other states, $\pi^{\beta}(\cdot;g)$ selects a random action from the set of available options provided by the environment. The pre-trained LLM is Llama-3-8B-Instruct with the following parameters: temperature: 1, top k: 40, maximum number of tokens: 200.
Each episode terminates either upon purchasing an object or reaching the step limit of 200. Using $\pi^{\beta}(\cdot;g)$, we collect 5,000 unlabeled trajectories spanning 1,500 unique instructions.
C.4 Step 1: Abstraction Function
The abstraction function utilizes contextual understanding of LLMs to identify goal-relevant features. The state abstraction operator is implemented as a collection of Python functions built on top of the feature selection made by a prompted language model, $F(\ \cdot\ ;g)=\mathrm{LLM}^{abstrct}(g)(\cdot)$. Using LLM powered Python functions instead of directly applying the LLM to create an abstraction of each state reduces the number of LLM calls from $|\mathcal{X}||G^{tr}|$ to $|G^{tr}|$ and ensures that the abstraction is consistent across all states. The prompt for generating the code includes contextual information about the environment, a list of features, and a description of the goal, instructing the LLM to create a function that removes features of a state that are irrelevant to achieving the specified goal. Additionally, our state abstraction functions separate the set of relevant features into two subsets: $s^{g}_{\phi}$ and $s^{g}_{\bar{\phi}}$, so that $s^{g}_{\phi}\cup s^{g}_{\bar{\phi}}$. The features in $s^{g}_{\phi}$ are the ones which are necessary for identifying if the underlying low-level state achieves $g$, i.e. if $x\in\phi(g)$. The remaining features, which are relevant for solving the task specified by $g$ but not strictly necessary for identifying if this goal is achieved are found in $s^{g}_{\bar{\phi}}$. We introduce this separation of abstracted features to even further reduce the dimensionality of the state space for hindsight labeling (see next section).
C.4.1 BabyAI adaptation
@@
To adopt the state abstraction function in for the BabyAI environment, we prompt an LLM with the given goal, a randomly sampled state representation as an example, and two in-context examples of the expected output. Figure C.4 shows the prompt template used. The LLM returns the goal relevant features which are then passed to a python function that processes states according to the following rules:
@@
In our experiments, we use the Llama-3-70B-Instruct language model with the following parameters: {temperature: 0, top k: 1, maximum number of tokens: 8000}.
C.4.2 Webshop adaptation
The WebShop environment includes predefined state abstractions that simplify the original HTML representation of the website. In our experiments, we utilize these abstractions, which are shared across all goals rather than being goal-specific, as in BabyAI. Furthermore, given the demonstrated ability of LLMs to parse HTML (Gur et al., 2023), we are confident that TEDUO Step 1 could replicate a similar abstraction. Since WebShop is non-Markovian, we ensure Markovian properties by using historical states–concatenating all states from the initial to the current state. References to ”states” in the following discussion refer to these history-augmented states.
@@ C.4.2 Webshop adaptation
Figure C.4: Prompt template for selecting the relevant features to achieve the goal in the BabyAI environment.
@@ 27. The agent is facing up
Figure C.5: An example of BabyAI textualized state before state abstraction.
@@
Figure C.6: Textualized state from C.5 after applying state abstraction for the goal “pick up a ball”.
C.5 Step 1: Reward Shaping
@@
In the first stage, a large language model LLM, here Llama-3-70B-Instruct, is utilized to generate a supervised dataset of labeled goals, $\{(s^{g},r^{g}):r^{g}=\mathrm{LLM}^{rwrd}(s^{g};g),s^{g}\in\tilde{{\mathcal{S}}}^{g}\}$ for each $g\in{\mathcal{G}}^{tr}$. We choose $\tilde{{\mathcal{S}}}^{g}$ as a small (up to 5000 states), diverse subset of the abstract space ${\mathcal{S}}^{g}$ and $r^{g}\in\{0,1\}$. The subset $\tilde{{\mathcal{S}}}^{g}$ is chosen so that for any two abstract states, the set of features relevant for goal-identifications is distinct, i.e. $\forall s^{g}_{1}\neq s^{g}_{2}\in\tilde{{\mathcal{S}}}^{g},s^{g}_{1,\phi}\neq s^{g}_{2,\phi}$. This maximizes the chances of including goal-states in $\tilde{{\mathcal{S}}}^{g}$, mitigating the potential issue of generating a highly-imbalanced dataset for training our proxy neural networks.
In the second stage, a collection of neural networks $R_{\theta}(\ \cdot\ ;g):{\mathcal{S}}^{g}\rightarrow\{0,1\}$ indexed by $g\in{\mathcal{G}}^{tr}$ is trained on the constructed supervised datasets.
C.5.1 BabyAI adaptation
In our experiments with the Baby AI environment, the state-labeling LLM is configured with parameters {temperature: 0, top-k: 1, max tokens: 8000}, and we use the prompt template shown in Figure C.7.
The state representations are transformed from text to a grid format. The network architecture consists of a small convolutional neural network with one convolutional layer (output dimension: 32, kernel size: (2,2)), followed by two linear layers (hidden dimension: 32, output dimension: 1). A Sigmoid activation function is applied after the final linear layer, and ReLU is used after all other layers. Dropout layers are added before each linear layer. The network is trained with the following hyperparameters: learning rate of 1e-5, maximum of 3000 epochs, and dropout rate of 0.1. The dataset is split into training and validation sets (90%/10%), and the model weights with the lowest validation loss are retained.
@@ C.5.1 BabyAI adaptation
Figure C.7: Prompt template for labeling states as goal states or not for the BabyAI environment.
C.5.2 Webshop adaptation
In TEDUO Step 1, goal-conditioned reward labeling is performed by prompting an LLM to evaluate the alignment between the instruction and the purchased product based on four criteria: category, attributes, options, and price. These criteria directly influence the true reward function defined by the environment. The resulting scores are combined using the formula specified in the WebShop environment to generate the synthetic reward. The prompt includes five example instruction evaluations (see Figure C.8). We use Llama-3.1-70B-Instruct with the following parameters: {temperature: 0, top-k: 1, max tokens: 100}. Since only terminal states (i.e., post-purchase states) require labeling in WebShop, the computational cost of this step is significantly reduced.
<|begin_of_text|><|start_header_id|>system<|end_header_id|>You are a helpful and honest judge of the fit between an instruction for purchasing an object and the proposed object. Your evaluation is based on whether the object meets four criteria: category, attributes, options, and price. Follow these steps: 1. Identify Attributes and Options - List the attributes mentioned in the instruction. - List the options selected in the instruction. 2.Evaluate Each Criterion - Category: Assign a value of 1 if the proposed objects product category matches the instruction; otherwise, assign 0. - Attributes: Count the number of attributes from the instruction that are correctly matched by the proposed object. - Options: Count the number of options selected by the user that are correctly matched by the proposed object. - Price: Assign a value of 1 if the proposed objects price is lower than or equal to the price specified in the instruction; otherwise, assign 0. <|eot_id|><|start_header_id|>user<|end_header_id|>{Example Instruction 1}}<|eot_id|> <|start_header_id|>assistant<|end_header_id|>{Example criteria 1}}<|eot_id|> <|start_header_id|>user<|end_header_id|>{Example Instruction 2}}<|eot_id|> <|start_header_id|>assistant<|end_header_id|>{Example critera 2}}<|eot_id|> <|start_header_id|>user<|end_header_id|>{Example Instruction 3}}<|eot_id|> <|start_header_id|>assistant<|end_header_id|>{Example criteria 3}}<|eot_id|> <|start_header_id|>user<|end_header_id|>{Example Instruction 4}}<|eot_id|> <|start_header_id|>assistant<|end_header_id|>{Example criteria 4}}<|eot_id|> <|start_header_id|>user<|end_header_id|>{Example Instruction 5}}<|eot_id|> <|start_header_id|>assistant<|end_header_id|>{Example criteria 5}}<|eot_id|> <|start_header_id|>user<|end_header_id|>{GOAL}}<|eot_id|> <|start_header_id|>assistant<|end_header_id|>
Figure C.8: Prompt template for synthetic reward in the Webshop environment.
C.6 Step 2: Offline Reinforcement Learning
C.6.1 BabyAI Adaptation
@@
$Q_{new}^{g}[s_{t},a_{t}]\leftarrow(1-\alpha)Q^{g}[s_{t},a_{t}]+\alpha\left(r^{g}[s_{t},a_{t}]+\gamma\max_{a}Q^{g}[s_{t+1},a]\right).$
In our experiments for each environment, the learning rate $\alpha$ is set to 0.1, and the discount factor $\gamma$ is set to 0.7. Subsequent states $s_{t+1}$ are restricted to transitions observed in $\mathcal{D}$. Iterations stop when $||Q_{new}^{g}-Q^{g}||_{\infty}<\epsilon$, where $\epsilon=1\times 10^{-6}$.
C.6.2 Webshop Adaptation
In the WebShop variant, given the deterministic nature of the environment and its unique starting state, we employ an improved filtered Behavioral Cloning method to solve the abstract MDPs. For each training goal $g$, we identify the terminal state with the highest reward $r^{g}$ among the collected samples. The corresponding trajectory is then refined by removing any potential loops. Finally, trajectories with a synthetic reward below a predefined threshold ($0.6$) are discarded.
C.7 Step 3: LLM Fine-tuning
Table C.8: Fine-tuning hyperparameters
Hyperparameter Value BabyAI Value Webshop
Batch size (per device) 10 4
Learning rate 2e-5 2e-6
Maximum Gradient norm 0.3 0.3
Warmup ratio 0.01 0.01
Maximum number of epochs 3 20
LORA rank 512 512
LORA alpha 512 512
LORA dropout 0.1 0.1
Split train/val ratio 0.1 0.1
Tensor type bf16 bf16
TEDUO’s step 3 involves fine-tuning a large language model using the generated supervised dataset ${\mathcal{D}}^{SFT}$. In this paper, the fine-tuned model is Llama-3-8B-Instruct. We use Low-Rank Adaptation ((Hu et al., 2021)) to reduce the compute cost. The hyperparameters used for the fine-tuning step are detailed in Table C.8. The model weights with the lowest validation loss are retained. The fine-tunings have been realized on a cluster of 4 A100 (80GB VRAM). The computing power provided in figure 5 is determined by multiplying the number of GPU hours by the peak Tflops (312 for A100 in bf16) and the estimated utilization rate (90%).
@@ C.7 Step 3: LLM Fine-tuning
Figure C.9: This prompt template is employed to generate a sequence of optimal actions to achieve the given goal while being in the given state for the BabyAI environment.
Figure C.10: This prompt template is employed to generate a sequence of optimal actions to achieve the given goal while being in the given state. It uses in-context learning and CoT prompting for the BabyAI environment.
C.8 Evaluation Setup
C.8.1 BabyAI adaptation
Goals. Training goals are defined as the goal contained in $\mathcal{G}^{tr}$, a subset of natural language instructions provided by BabyAI without any modifications. Testing goals differ both grammatically and semantically from training goals. They are derived from BabyAI’s original instructions, distinct from $\mathcal{G}^{tr}$, and reformulated using alternative phrasings and synonyms. Tables C.9, C.10, and C.11 provide the alternative formulations and synonyms for objects and colors used in these reformulations.
Baselines:
LLMs (vanilla). The vanilla Large Language Model baseline utilizes Llama-3-8B-Instruct or Llama-3-70B-Instruct prompted with the template shown in Figure C.9. This prompt provides basic information about the environment, current goal, and a textual (non abstracted) representation of the state.
LLMs (in-context + CoT). This baseline extends the vanilla LLM approach by using the prompt in Figure C.10, which includes detailed environment information, similar to a game manual, as described in (Wu et al., 2023). It also integrates expert demonstrations using textual grid examples and goals with their optimal action sequences. The CoT prompting technique is employed to guide the LLM through multi-step reasoning and self-reflection.
@@ C.8.1 BabyAI adaptation
C.8.2 Webshop adaptation
Environment. Since the WebShop environment does not provide multiple configurations, our evaluation focuses on generalization to new goals.
Goals. The training goals $\mathcal{G}^{tr}$ are a subset of the natural language instructions provided by the WebShop environment, used without modification. Similarly, the testing goals form a distinct subset of instructions generated using different seeds while ensuring no overlap with $\mathcal{G}^{tr}$. Given the grammatical and semantic diversity of the instructions, we do not perform any goal reformulation.
Baselines:
ReAct. This baseline leverages general-purpose LLMs with the ReAct prompting technique (Yao et al., 2023b), which is state-of-the-art for low-data settings in this environment. ReAct enhances reasoning in LLMs by integrating reasoning traces with action steps within the same prompt, enabling the model to reason through complex problems while simultaneously retrieving and validating relevant information. This feedback loop improves both accuracy and coherence. Vanilla prompting was excluded, as it failed to reach the purchasing step within the step limit (15), preventing it from achieving valid rewards.
Table C.9: Alternative formulations for the natural language commands.
@@ C.8.2 Webshop adaptation
Table C.10: Synonyms used for the objects.
@@
Table C.11: Synonyms used for the colors.